# Cosmic Ray Superdiffusion and Mirror Diffusion in Partially Ionized and Turbulent Medium

### v1

@@ Abstract
Understanding cosmic ray (CR) diffusion in a partially ionized medium is both crucial and challenging. In this study, we investigate CR superdiffusion in turbulent, partially ionized media using high-resolution 3D two-fluid simulations that treat ions and neutrals separately. We examine the influence of neutral-ion decoupling and the associated damping of turbulence on CR propagation in both transonic and supersonic conditions. Our simulations demonstrate that neutral-ion decoupling significantly damps velocity, density, and magnetic field fluctuations at small scales, producing spectral slopes steeper than those of Kolmogorov and Burgers scaling. We also identify an intermediate coupling regime in which neutrals remain partially coupled with ions, leading to a steepening of kinetic energy spectra in both fluids. Shocks can enhance the neutral-ion coupling, thereby reducing the differences between neutral and ion density structures. Moreover, the damping of magnetic field fluctuations substantially decreases pitch-angle scattering, which increases CR parallel mean free paths. As a result, CR perpendicular transport transitions between two distinct superdiffusive regimes: a scattering-dominated regime with perpendicular displacement proportional to $t^{3/4}$, and a scattering-free regime dominated by magnetic field line wandering, with perpendicular displacement scaling as $t^{3/2}$. When the pitch angle is large, the effects of magnetic mirroring—naturally arising in magnetohydrodynamic turbulence—become significant, enhancing the confinement (but not fully trapping) of CRs and inducing oscillations in their perpendicular displacement. These results highlight the necessity of incorporating two-fluid effects for accurately modeling CR transport in partially ionized environments such as molecular clouds and dense interstellar clumps.
@@ 1 Introduction
A key determinant of CR transport is the nature of turbulent magnetic fields. Fluctuating fields introduce pitch-angle scattering and govern the effective diffusion of CRs along field lines (Jokipii, 1966; Jokipii & Parker, 1969; Qin et al., 2002; Lazarian & Xu, 2021). A growing body of work has revealed that CR transport can deviate significantly from the classical diffusion paradigm in the presence of magnetohydrodynamic (MHD) turbulence (Comisso & Sironi, 2018, 2019; Beattie et al., 2022; Lazarian et al., 2023; Zhang & Xu, 2024; Marcowith, 2025). In particular, turbulent magnetic field wandering (Lazarian & Vishniac (1999)) induces superdiffusive transport perpendicular to mean magnetic field (e.g., Yan & Lazarian, 2008; Xu & Yan, 2013; Lazarian & Yan, 2014; Hu et al., 2022), which in combination with the effects of CR bouncing in magnetic mirrors (Cesarsky & Kulsrud, 1973) results in parallel propagation that strongly depends on the CR pitch angle, i.e., in the process termed ”mirror diffusion” (Lazarian & Xu, 2021; Xu & Li, 2023). However, many studies assume a fully ionized plasma or utilize single-fluid approximations. In dense clumps, on the contrary, the gas is cold ($\sim$10–100 K) and only weakly ionized, resulting in additional physical processes such as ion-neutral collisions, charge exchange and turbulence damping (e.g., Draine, 1986; Balsara, 1996; Tilley & Balsara, 2011; Xu et al., 2015). These effects can alter both the local turbulence spectrum and CR scattering rates (Tilley & Balsara, 2010; McKee et al., 2010; Burkhart et al., 2015; Hu et al., 2024). Consequently, a two-fluid (ion + neutral) description is critical to fully capture the microphysics that governs CR propagation in these environments. By incorporating two-fluid effects, one can better connect observations—such as locally enhanced gamma-ray emission in molecular clouds—to the underlying plasma and magnetic field conditions. Consequently, a two-fluid (ion + neutral) description is critical to fully capture the microphysics that governs CR propagation in these environments. Incorporating two-fluid effects can better connect observations, such as locally enhanced gamma-ray emission in molecular clouds, to the underlying plasma and magnetic field conditions.
@@
Despite the recognized importance of two-fluid processes, direct numerical simulations of this regime remain challenging because the need to resolve large Alfvén speeds in the neutral-ion decoupling layers quickly inflates computational costs. Earlier work has shown that neutral-ion decoupling can significantly modify the turbulent energy spectra of ions and neutrals (Tilley & Balsara, 2010; McKee et al., 2010; Burkhart et al., 2015). In particular, Hu et al. (2024) found that large density fluctuations in ions and neutrals can drive substantial spatial variations in the ionization fraction and local neutral-ion decoupling scale, leading to a globally intermediately coupled status. In this intermediately coupled regime, turbulence in both fluids can be damped, resulting in steeper kinetic energy spectra than the familiar Kolmogorov profile, which profoundly influences CR propagation.
@@
In this work, we extend existing theories of CR transport and magnetic turbulence by incorporating essential two-fluid physics of neutral-ion decoupling and turbulence damping. We perform two-fluid MHD simulations using the AthenaK code (Stone et al., 2024), varying the ionization fraction and coupling strength to explore different regimes of ion-neutral interaction. By examining the resulting CR scattering properties and superdiffusive behavior, we aim to clarify how partial ionization modifies turbulence and, consequently, CR transport in dense interstellar environments.
@@
This paper is organized as follows. § 2 briefly review the theoretical consideration of CR superdiffusion in a magnetized and partially ionized medium. § 3 describes the 3D two-fluid simulations and numerical approach of modeling CR propagation used in this study. In § 4, we analyze the statistics of turbulence and CR diffusion in the partially ionized medium across different neutral-ion coupling statuses. We conclude with a summary of our findings in § 5.
@@ 2 Theoretical consideration
2.1 Decoupling of neutrals and ions
The coupling between ions and neutrals is commonly characterized by two collisional frequencies. The neutral–ion collision frequency is defined as $\nu_{ni}=\gamma_{\rm d}\rho_{i}=\gamma_{\rm d}\xi_{i}(\rho_{i}+\rho_{n})$ while the ion–neutral collision frequency is given by $\nu_{in}=\gamma_{\rm d}\rho_{n}$, respectively (Shu, 1992). Decoupling of neutrals from ions begins when the rate at which turbulent energy cascades becomes comparable to $\nu_{ni}$. Because $\nu_{in}$ greatly exceeds $\nu_{ni}$ in weakly ionized media, ions separate from neutrals on much smaller scales than those marking the decoupling of neutrals; hence, our focus here is primarily on the neutral–ion decoupling process.
@@ 2.1 Decoupling of neutrals and ions
Although MHD turbulence was once approximated as a superposition of linear waves (Giacalone & Jokipii, 1999), it is now recognized as an inherently non-linear process. In particular, within the strong turbulence regime under a critical balance condition, any wavelike behavior of Alfv’en is short-lived, persisting for less than a wave period (Goldreich & Sridhar, 1995; Lazarian & Vishniac, 1999). With the dynamically coupled turbulent mixing motion in the perpendicular direction and the wave-like motion in the parallel direction, the scale-dependent anisotropy of MHD turbulence develops. For Alfvénic turbulence  [^1] For a more in-depth discussion on neutral-ion decoupling scales of the three MHD modes (Alfvén, fast, and slow), see Xu et al. (2015, 2016)., which typically dominates the turbulent energy budget (Cho et al., 2002; Hu et al., 2021), this anisotropy implies that the decoupling length is itself anisotropic—the parallel component can exceed the perpendicular one, so that the latter is smaller than earlier estimates (Xu et al., 2015, 2016; Hu et al., 2024). The perpendicular decoupling scale in strong MHD turbulence regimes is expressed as (Xu et al., 2015):
$k_{{\rm dec},\bot}$ =$=\begin{cases}\nu_{\rm ni}^{3/2}L_{\rm inj}^{1/2}v_{\rm inj}^{-3/2},~{}~{}~{}M_{A}>1\\ \nu_{\rm ni}^{3/2}L_{\rm inj}^{1/2}v_{\rm inj}^{-3/2}M_{A}^{-1/2},~{}~{}~{}M_{A}<1\\ \end{cases},$ (2)
@@
Hu et al. (2024) further revealed that, in a turbulent MHD environment, large density fluctuations in both ions and neutrals lead to variations in the Alfvén speed and ionization fraction. These variations cause the parallel and perpendicular neutral–ion decoupling scales to span over an extended range. Consequently, the overall coupling status between ions and neutrals can be categorized into three regimes: (i) a strongly coupled regime, where scales larger than the maximum decoupling length engender single-fluid behavior; (ii) a intermediately coupled regime, corresponding to intermediate scales where the neutral and ions can be either locally coupled or locally decoupled, , depending on the local physical conditions; and (iii) a fully decoupled regime, at scales below the minimum decoupling scale, in which neutrals evolve an independent hydrodynamic cascade while ions follow a separate MHD cascade if turbulent energy is injected at these small scales.
@@ 2.2 Neutral-ion collisional damping of MHD turbulence
Strongly coupled regime: In the strongly coupled regime, ions and neutrals move in unison, effectively forming a single fluid that transports the MHD turbulence. Both species share and sustain the cascading turbulent energy without any noticeable separation in their dynamics.
@@
Intermediate coupling: Transitioning to the intermediately coupled regime, neutrals begin to decouple locally from the ions. Although ions continue to experience frequent collisions with nearby neutrals—extending down to the ion–neutral decoupling scale—the local decoupling of neutrals leads to a significant modification of the ion dynamics. The MHD turbulence in ions is strongly damped by the frictional forces arising from these collisions. In this regime, the hydrodynamic turbulence in neutrals is also modified and damped due to interactions with ions (Hu et al., 2024).
@@
Fully decoupled regime: In the fully decoupled regime, the behavior of the two components diverges markedly. Ions remain subject to frequent collisions with the surrounding neutrals, which results in the damping of their MHD turbulent cascade. Conversely, neutrals develop an independent hydrodynamic cascade. This decoupling leads to a clear separation in the dynamical evolution: while the ion turbulence remains damped by collisional friction, the neutrals evolve under a purely hydrodynamic regime.
@@
The damping of MHD turbulence leads to a marked reduction in fluctuations of density, velocity, and especially the magnetic field. This diminished fluctuations in the magnetic field curtail the efficiency of pitch angle scattering for CRs, thereby altering their transport properties along the magnetic field lines.
Figure 1: An example of the scattering diffusion and mirror diffusion. A particle with a Larmor radius of 10 cells and an initial pitch angle cosine of 0.25 is ejected into the simulation A4 ($M_{s}=1.24$ and $M_{A}=1.21$). The particle trajectory is plotted in the 3D volume rendering cube of normalized magnetic field strength. The trajectory is colored to show the particle’s evolution time: time increases from blue to red. In the regimes dominated by a magnetic field mirror, the particle is oscillating back and forth but gradually diffusion out of the mirror.
2.3 CRs’ perpendicular superdiffusion induced by wandering magnetic field lines
CRs’ propagation along the magnetic field lines can be separated into two regimes based on the relationship between their mean free path $\langle\lambda_{\parallel}\rangle$ (parallel to the magnetic field).
(1) When $\langle\lambda_{\parallel}\rangle>L_{\rm inj}$: the scattering of CRs is negligible, and they move along magnetic field lines in a ballistic manner. The divergence of magnetic field lines primarily determines their diffusion perpendicular to the mean magnetic field. Specifically, the magnetic field lines separate superdiffusively in space Lazarian & Vishniac (1999):
$\langle z^{2}\rangle\sim\frac{x^{3}}{27L_{\rm inj}}M_{A}^{4},~{}~{}{M_{A}\leq 1},$ (3)
$\langle z^{2}\rangle\sim\frac{x^{3}}{27L_{\rm inj}}M_{A}^{3},~{}~{}{M_{A}>1},$
where $x$ is the distance traced along magnetic field lines that are initially infinitely close to each other [^2] A peculiar feature of this law is that for sufficiently large $x$, the line separation can be independent of the initial line separation as discussed in Eyink et al. (2011)., while $\sqrt{\langle z^{2}\rangle}$ is the averaged separation of magnetic field lines perpendicular direction gained as the result of this tracing. Note that the dependence on $M_{A}$ is different for sub-Alfvénic and super-Alfvénic turbulence.
The magnetic field wandering imprints its law on the CRs moving along magnetic field lines (Xu & Yan, 2013). Those exhibit superdiffusion in the direction perpendicular to the magnetic field (Yan & Lazarian, 2008; Lazarian & Yan, 2014). This superdiffusion in space is broadly related to the Richardson dispersion of magnetic field lines in time Eyink et al. (2013), but is not equivalent to it. The relation between the two can be exemplified by considering CR moving ballistically along the wandering magnetic field lines with constant velocity. After time $t$, the dispersion of CR positions is:
$\langle z^{2}\rangle\sim\frac{t^{3}}{27L_{\rm inj}}M_{A}^{4},~{}~{}{M_{A}\leq 1},$ (4)
$\langle z^{2}\rangle\sim\frac{t^{3}}{27L_{\rm inj}}M_{A}^{3},~{}~{}{M_{A}>1},$
The functional dependence of $\sqrt{\langle z^{2}\rangle}$ on $t$ formally coincides with the law of the Richardson dispersion of magnetic field lines. However, the process described above is not the advection of magnetic field lines by turbulence described by Richardson diffusion. Indeed, in our example above, the magnetic field lines are assumed to be static.
Note that Eq. 4 is valid for strong turbulence, specifically for scales smaller than $l_{\rm tr}=L_{\rm inj}M_{A}^{2}$ when $M_{A}\leq 1$ or $l_{A}=L_{\rm inj}M_{A}^{-3}$ when $M_{A}>1$.  [^3] In the weak turbulence regime (i.e., scales larger than $l_{\rm tr}$ or $l_{A}$), turbulence is wave-like rather than anisotropic-eddy-like. In this regime, the separation of magnetic field lines follows a random walk, resulting in normal diffusion of CRs: $\langle z^{2}\rangle\sim tL_{\rm inj}M_{A}^{4},~{}~{}{M_{A}\leq 1},$ (5) $\langle z^{2}\rangle\sim tL_{\rm inj}M_{A}^{-3},~{}~{}{M_{A}>1}.$ For a detailed derivation of these equations, see Yan & Lazarian (2008); Lazarian & Yan (2014); Hu et al. (2022).
@@ 2.3 CRs’ perpendicular superdiffusion induced by wandering magnetic field lines
Table 1: The setup of two-fluid simulations at a resolution of $1200^{3}$ is summarized here. The sonic Mach number ($M_{s}$) and Alfvénic Mach number ($M_{A}$) represent the instantaneous RMS values at each snapshot. The plasma compressibility is given by $\beta=2(M_{A}/M_{s})^{2}$. The ratio of the neutral-ion collisional frequency to the eddy turnover rate is expressed as $(\gamma_{\rm d}\rho_{i})/(v_{\rm inj}/L_{\rm inj})$, which characterizes the coupling strength. The theoretically expected parallel and perpendicular components of the neutral-ion decoupling wavenumber are denoted as $k_{\rm dec,\parallel}$ and $k_{\rm dec,\bot}$, respectively, while the corresponding neutral-ion decoupling scales are given by $l_{\rm dec,\parallel}$ and $l_{\rm dec,\bot}$. The numerical dissipation scale is represented by $l_{\rm dis}$.

### v2

# Cosmic Ray Perpendicular Superdiffusion and Parallel Mirror Diffusion in a Partially Ionized and Turbulent Medium

@@ Abstract
Understanding cosmic ray (CR) diffusion in a partially ionized medium is both crucial and challenging. In this study, we investigate CR perpendicular superdiffusion and parallel transport in turbulent, partially ionized media using high-resolution 3D two-fluid simulations that treat ions and neutrals separately. We examine the influence of neutral-ion decoupling and the associated damping of turbulence on CR propagation in both transonic and supersonic conditions. Our simulations demonstrate that neutral-ion decoupling significantly damps velocity and magnetic field fluctuations at small scales, producing spectral slopes steeper than those of Kolmogorov and Burgers scaling. In supersonic turbulence, large-scale shock motion is not subject to damping and generates small-scale density enhancements. Moreover, the damping of magnetic field fluctuations substantially decreases pitch-angle scattering, which, however, only slightly affects the CR parallel mean free path $\lambda_{\|}$, due to the nonresonant mirror interactions of CRs. In the direction perpendicular to the mean magnetic field, we identify two regimes of the perpendicular superdiffusion of CRs: a diffusive regime ($\lambda_{\|}<L_{\rm inj}$, where $L_{\rm inj}$ is turbulence injection scale) with perpendicular separation of CR proportional to $t^{3/4}$, and a ballistic-like regime ($\lambda_{\|}>L_{\rm inj}$), with perpendicular separation scaling as $t^{3/2}$. At initially large pitch angles, the effects of magnetic mirroring—naturally arising in magnetohydrodynamic turbulence—become significant, enhancing the confinement of CRs and resulting in $\lambda_{\|}<L_{\rm inj}$, despite the damping effect. These results imply that large-pitch-angle CRs can be well confined in the cold ISM, such as molecular clouds.
@@ 1 Introduction
A key determinant of CR transport is the nature of turbulent magnetic fields. Fluctuating fields introduce pitch-angle scattering and magnetic mirroring, which can affect the diffusion of CRs along field lines (Jokipii, 1966; Jokipii & Parker, 1969; Qin et al., 2002; Lazarian & Xu, 2021). A growing body of work has revealed that CR transport can deviate significantly from the classical diffusion paradigm in the presence of magnetohydrodynamic (MHD) turbulence (Comisso & Sironi, 2018, 2019; Hopkins et al., 2022b; Kempski & Quataert, 2022; Beattie et al., 2022; Lazarian et al., 2023; Lemoine, 2023; Kempski et al., 2023; Zhang & Xu, 2024; Marcowith, 2025). In particular, turbulent magnetic field lines’ superdiffusion (Lazarian & Vishniac, 1999) induces superdiffusive transport perpendicular to mean magnetic field (e.g., Yan & Lazarian, 2008; Xu & Yan, 2013; Lazarian & Yan, 2014; Hu et al., 2022). Lazarian & Xu (2021) combine the effects of CR bouncing in magnetic mirrors (Cesarsky & Kulsrud, 1973) results in parallel propagation that strongly depends on the CR pitch angle, i.e., in the process termed ”mirror diffusion” (Lazarian & Xu, 2021; Xu & Li, 2023; Barreto-Mota et al., 2025; Xiao et al., 2025). However, most earlier studies assume a fully ionized plasma or adopt single-fluid approximations. In molecular cloud, on the contrary, the gas is cold ($\sim$10–100 K) and only weakly ionized, resulting in additional physical processes such as ion-neutral collisions and ion-neutral collisional damping of turbulence (e.g., Draine, 1986; Balsara, 1996; Tilley & Balsara, 2011; Xu et al., 2015). These effects can alter both the local turbulence spectrum and CR scattering rates (Tilley & Balsara, 2010; McKee et al., 2010; Burkhart et al., 2015; Hu et al., 2024). Consequently, a two-fluid (ion + neutral) description is critical to fully capture the physics that governs CR propagation in these environments. By incorporating two-fluid effects, one can better connect observations—such as locally enhanced gamma-ray emission in molecular clouds—to the underlying plasma and magnetic field conditions.
@@
Despite the high numerical cost to resolve extremely large Alfvén speeds in the neutral-ion decoupled status, the two-fluid approach is necessary, as earlier work has shown that neutral-ion decoupling can significantly modify the turbulent energy spectra of ions and neutrals (Tilley & Balsara, 2010; McKee et al., 2010; Burkhart et al., 2015). In particular, Hu et al. (2024) found that large density fluctuations in ions and neutrals can drive substantial spatial variations in the ionization fraction and local neutral-ion coupling status. Before reaching the locally smallest decoupling scale, i.e., the fully decoupled regime, neutrals can remain partially coupled with ions in an intermediate coupled regime. In the intermediately coupled and fully decoupled regimes, turbulence can be damped, resulting in steeper kinetic and magnetic field spectra than the familiar Kolmogorov profile, which profoundly influences CR propagation.
@@
In this work, we extend existing theories of magnetic turbulence and CR transport by incorporating essential two-fluid physics of neutral-ion decoupling and the associated collisional damping of turbulence. We perform two-fluid MHD simulations using the AthenaK code (Stone et al., 2024), varying the ionization fraction and coupling strength to explore different regimes of ion-neutral interaction. By examining the resulting CR transport properties, we aim to clarify how partial ionization modifies turbulence and, consequently, CR transport in dense interstellar environments.
@@
This paper is organized as follows. § 2 provides a brief review of the theoretical framework for CR perpendicular superdiffusion and parallel mirror diffusion in a magnetized, partially ionized medium. § 3 describes the 3D two-fluid simulations and numerical approach of modeling CR propagation used in this study. In § 4, we analyze the statistics of turbulence and CR diffusion in the partially ionized medium at different degrees of neutral-ion coupling. We conclude with a summary of our main findings in § 5.
@@ 2 Theoretical consideration
2.1 Decoupling of neutrals from ions
The coupling between ions and neutrals is commonly characterized by two collisional frequencies. The neutral–ion collision frequency is defined as $\nu_{ni}=\gamma_{\rm d}\rho_{i}=\gamma_{\rm d}\xi_{i}(\rho_{i}+\rho_{n})$ while the ion–neutral collision frequency is given by $\nu_{in}=\gamma_{\rm d}\rho_{n}$, respectively (Shu, 1992). Here $\gamma_{\rm d}$ represents the drag coefficient that couples ions and neutrals through the collisional exchange of momentum, while $\xi_{i}=\rho_{i}/(\rho_{i}+\rho_{n})$ is the ionization fraction. $\rho_{i}$ and $\rho_{n}$ are ion mass density and neutral mass density, respectively. Decoupling of neutrals from ions begins when the rate at which turbulent energy cascades becomes comparable to $\nu_{ni}$. Because $\nu_{in}$ greatly exceeds $\nu_{ni}$ in weakly ionized media, ions separate from neutrals on much smaller scales than those marking the decoupling of neutrals; hence, our focus here is primarily on the neutral–ion decoupling process.
@@ 2.1 Decoupling of neutrals from ions
Although MHD turbulence was approximated as a superposition of linear waves in some earlier studies (Giacalone & Jokipii, 1999), it is now recognized as an inherently non-linear process. In particular, within the strong turbulence regime under a critical balance condition, any Alfvén wave behavior is short-lived, persisting for a wave period (Goldreich & Sridhar, 1995; Lazarian & Vishniac, 1999). With the dynamically coupled turbulent mixing motion in the perpendicular direction and the wave-like motion in the parallel direction, the cascade of the MHD turbulence is mainly along the perpendicular direction, and the scale-dependent anisotropy of MHD turbulence develops (see Xu 2019 for a review). For Alfvénic turbulence  [^1] For a more in-depth discussion on neutral-ion decoupling scales of the three MHD modes (Alfvén, fast, and slow), see Xu et al. (2015, 2016)., which typically dominates the turbulent energy budget (Cho et al., 2002; Hu et al., 2021), this anisotropy implies that the decoupling length is itself anisotropic—the parallel component can exceed the perpendicular one, so that the latter is smaller than earlier estimates (Xu et al., 2015, 2016; Hu et al., 2024). The perpendicular decoupling scale in strong MHD turbulence regimes is expressed as (Xu et al., 2015):
[…]
@@
Figure 1: An example of the scattering diffusion and mirror diffusion. A particle with a Larmor radius of 10 cells and an initial pitch angle cosine of 0.25 is ejected into the simulation A4 ($M_{s}=1.24$ and $M_{A}=1.21$). The particle trajectory is plotted in the 3D volume rendering cube of normalized magnetic field strength. Panel (a): The trajectory is colored to show the time, increasing from blue to red. Panel (b): The trajectory is colored to show $|\mu|$, where $\mu$ is the cosine pitch angle, increasing from dark to yellow. Magnetic field lines are colored to show the normalized magnetic field strength. In the regimes dominated by a magnetic field mirror, the particle is reflected back and forth by multiple mirrors, while undergoing superdiffusion perpendicular to the mean magnetic field. It can reenter the scattering diffusion regime due to the gradual stochastic change of pitch angle.
@@ 2.2 Neutral-ion collisional damping of MHD turbulence
Hu et al. (2024) further revealed that, in compressible MHD turbulence, large density fluctuations in both ions and neutrals lead to spatial variations in the Alfvén speed and ionization fraction. These variations cause the parallel and perpendicular neutral–ion decoupling scales to span over an extended range. Because of the spatial variation of the decoupling scales, the overall coupling status of neutrals with ions can be categorized into three regimes.
@@
Strongly coupled regime: In the strongly coupled regime, where scales are larger than the maximum parallel decoupling length, i.e., $k<k_{\rm dec,\|}$, ions and neutrals move in unison, effectively forming a single fluid that carries the MHD turbulence. Both species share and sustain the cascading turbulent energy without any noticeable separation in their dynamics.
@@
Intermediate coupling: The intermediately coupled regime corresponds to intermediate scales where the neutral and ions can be either locally coupled or locally decoupled, depending on the local physical conditions. Although ions continue to experience frequent collisions with nearby neutrals—extending down to the ion–neutral decoupling scale—the local decoupling of neutrals leads to a significant modification of the ion dynamics. The MHD turbulence in ions is strongly damped by the friction forces arising from these collisions. In this regime, the hydrodynamic turbulence in neutrals is also modified locally (Hu et al., 2024).
@@
Fully decoupled regime: In the fully decoupled regime, where the scales are below the minimum perpendicular decoupling scale, i.e., $k>k_{\rm dec,\bot}$, the behavior of the two components diverges markedly. Ions remain subject to frequent collisions with the surrounding neutrals, which results in the damping of their MHD turbulent cascade. Conversely, as neutrals fully decouple from ions, neutrals develop an independent hydrodynamic cascade. This decoupling of neutrals from ions leads to a clear separation in the dynamical evolution: while the ion turbulence remains damped by collisional friction, the neutrals evolve under a purely hydrodynamic regime.
@@
The damping of MHD turbulence leads to a marked reduction in fluctuations of density, velocity, and especially the magnetic field. These diminished fluctuations in the magnetic field curtail the efficiency of pitch angle scattering for CRs and also affect the properties of magnetic mirrors, thereby altering CR’s transport.
@@
Table 1: The setup of two-fluid simulations at a resolution of $1200^{3}$ is summarized here. The sonic Mach number ($M_{s}$) and Alfvénic Mach number ($M_{A}$) represent the instantaneous RMS values at each snapshot. The plasma compressibility is given by $\beta=2(M_{A}/M_{s})^{2}$. The ratio of the neutral-ion collisional frequency to the eddy turnover rate at injection scale is expressed as $(\gamma_{\rm d}\rho_{i})/(v_{\rm inj}/L_{\rm inj})$, which characterizes the coupling strength at injection scale. The theoretically expected parallel and perpendicular components of the neutral-ion decoupling wavenumber are denoted as $k_{\rm dec,\parallel}$ and $k_{\rm dec,\bot}$, respectively, while the corresponding neutral-ion decoupling scales are given by $l_{\rm dec,\parallel}$ and $l_{\rm dec,\bot}$. The numerical dissipation scale is represented by $l_{\rm dis}$.
2.3 CRs’ perpendicular superdiffusion induced by perpendicular superdiffusion of magnetic field lines
The perpendicular superdiffusion of CRs depends on their parallel diffusion behavior. It can be separated into two regimes based on the relationship between their mean free path $\lambda_{\parallel}$ (parallel to the magnetic field) and the injection scale of turbulence $L_{\rm inj}$.
(1) When $\lambda_{\parallel}>L_{\rm inj}$: the scattering and mirroring of CRs are negligible, and they move along magnetic field lines in a ballistic manner. The divergence of magnetic field lines primarily determines their diffusion perpendicular to the mean magnetic field. Specifically, the magnetic field lines separate superdiffusively in space (Lazarian & Vishniac, 1999):
$\langle z^{2}\rangle\sim\frac{x^{3}}{27L_{\rm inj}}M_{A}^{4},~~{M_{A}\leq 1},$ (3)
$\langle z^{2}\rangle\sim\frac{x^{3}}{27L_{\rm inj}}M_{A}^{3},~~{M_{A}>1},$
where $x$ is the distance traced along magnetic field lines that are initially close to each other [^2] A peculiar feature of this law is that for sufficiently large $x$, the line separation can be independent of the initial line separation as discussed in Eyink et al. (2011). but larger than the damping scale, while $\sqrt{\langle z^{2}\rangle}$ is the averaged separation of magnetic field lines in the perpendicular direction gained as the result of this tracing. Note that the dependence on $M_{A}$ is different for sub-Alfvénic and super-Alfvénic turbulence.
The magnetic field superdiffusion imprints its law on the CRs moving along magnetic field lines (Xu & Yan, 2013). Those exhibit superdiffusion in the direction perpendicular to the mean magnetic field (Yan & Lazarian, 2008; Lazarian & Yan, 2014). This superdiffusion in space is broadly related to the Richardson dispersion of magnetic field lines in time Eyink et al. (2013), but is not equivalent to it. The relationship between the two can be exemplified by considering CR moving ballistically along the magnetic field lines with a negligible change in pitch angle and thus a constant parallel velocity. After time $t$, the CR traveled a distance of $x\sim c\mu t$ along the magnetic field line, where $c$ is the speed of light and $\mu$ is the pitch angle cosine. The CR’s perpendicular separation can be expressed as:
$\langle z^{2}\rangle\sim\frac{(c\mu t)^{3}}{27L_{\rm inj}}M_{A}^{4},~~{M_{A}\leq 1},$ (4)
$\langle z^{2}\rangle\sim\frac{(c\mu t)^{3}}{27L_{\rm inj}}M_{A}^{3},~~{M_{A}>1},$
in which the functional dependence of $\sqrt{\langle z^{2}\rangle}$ on $t$ formally coincides with the law of the Richardson dispersion of magnetic field lines. However, the process described above is not the advection of magnetic field lines by turbulence described by Richardson diffusion, but the stochasticity of turbulent field lines due to their turbulent reconnection (Lazarian & Vishniac, 1999). Indeed, in our example above, the magnetic field lines are assumed to be static.
Note that Eq. 4 is valid for strong turbulence, specifically for scales smaller than $l_{\rm tr}=L_{\rm inj}M_{A}^{2}$ when $M_{A}\leq 1$ or $l_{A}=L_{\rm inj}M_{A}^{-3}$ when $M_{A}>1$. In the weak turbulence regime (i.e., scales larger than $l_{\rm tr}$ or $l_{A}$), turbulence is hydrodynamic-like rather than anisotropic-eddy-like. In this regime, the separation of magnetic field lines follows a random walk, and the CRs’ perpendicular separation can be expressed as:
$\langle z^{2}\rangle\sim\frac{c\mu tL_{\rm inj}}{3}M_{A}^{4},~~{M_{A}\leq 1},$ (5)
$\langle z^{2}\rangle\sim\frac{c\mu tL_{\rm inj}}{3}M_{A}^{-3},~~{M_{A}>1},$
which suggests a normal perpendicular diffusion of CRs at large scales. For a detailed derivation of these equations, see Yan & Lazarian (2008); Lazarian & Yan (2014); Hu et al. (2022). The transition from small-scale ($L_{\rm inj}$) perpendicular superdiffusion to large-scale ($L_{\rm inj}$) perpendicular normal diffusion has been numerically tested by Zhang & Xu (2023).